### v1

@@ 4 A priori estimates and stability for BSDEs
In this section, we establish the stability of the mean-field BSDE (2.5), as stated in Theorem 4.6. The strategy of the proof is as follows. We first derive a priori upper bounds on the $\mathbb{S}^{\infty}$ norm of $Y$ and the BMO norm of $\int_{0}^{\cdot}Z_{s}\cdot dW_{s}$ for any solution $(Y,Z)$ to the BSDE (2.5). These estimates, together with fundamental properties of BMO martingales, lead to the stability result in Theorem 4.6. We begin with a useful estimate that will play an important role in the analysis.
@@
Suppose that (H0) and (Hq) hold. Let $(Y,Z)\in\mathbb{S}^{\infty}\times\mathbb{H}^{2}$ be a bounded solution to the mean-field BSDE (2.5). Then the stochastic integral process $\left(\int_{0}^{t}Z_{s}\cdot dW_{s}\right)_{0\leq t\leq T}$ defines a BMO martingale. Moreover, its BMO norm satisfies the following bound:
@@
The argument follows the two steps scheme of (briand2008stobsdes, p. 831) with modified $C^{2}$ function
$\varphi(x):=\frac{e^{3K_{z}|x|}-1-6cK_{z}|x|}{(3K_{z})^{2}},\qquad x\in\mathbb{R}.$
@@
Step 1 (Estimate of $\|Z\|_{\mathrm{BMO}}$ in terms of $\|Z\|_{\mathbb{H}^{2}}$). Fix $c=\tfrac{1}{2}$ then $\varphi(0)=0$ and, for all $x\in\mathbb{R}$,
$\varphi^{\prime}(x)=\frac{e^{3K_{z}|x|}-1}{3K_{z}}\operatorname{sgn}(x),\qquad\varphi^{\prime\prime}(x)=e^{3K_{z}|x|},\qquad\varphi^{\prime\prime}(x)-3K_{z}\,|\varphi^{\prime}(x)|\equiv 1.$
@@
Let $\tau$ be any stopping time and apply Itô’s formula to $\varphi(Y)$ and using the growth estimate 4.8 for $g$ together with the symmetry $\widetilde{\mathbb{E}}|\tilde{Z}_{s}|^{2}=\mathbb{E}|Z_{s}|^{2}$ yields
$\varphi(Y_{\tau})$ $=\varphi(Y_{T})+\!\int_{\tau}^{T}\!\Big(-\varphi^{\prime}(Y_{s})\,\widetilde{\mathbb{E}}[g(s,Z_{s},\tilde{Z}_{s})]-\tfrac{1}{2}\,\varphi^{\prime\prime}(Y_{s})|Z_{s}|^{2}\Big)ds-\!\int_{\tau}^{T}\!\varphi^{\prime}(Y_{s})Z_{s}\cdot dW_{s}$
$\leq\varphi(Y_{T})+\!\int_{\tau}^{T}\!\Big(\tfrac{3}{2}|\varphi^{\prime}(Y_{s})|-\tfrac{1}{2}\varphi^{\prime\prime}(Y_{s})\Big)|Z_{s}|^{2}ds+\!\int_{\tau}^{T}\!|\varphi^{\prime}(Y_{s})|\Big(K+\tfrac{3}{2}K_{\tilde{z}}\,\mathbb{E}|Z_{s}|^{2}\Big)ds$
$\leq\varphi(Y_{T})-\tfrac{1}{2}\!\int_{\tau}^{T}\!|Z_{s}|^{2}ds+\!\int_{\tau}^{T}\!|\varphi^{\prime}(Y_{s})|\Big(K+\tfrac{3}{2}K_{\tilde{z}}\,\mathbb{E}|Z_{s}|^{2}\Big)ds-\!\int_{\tau}^{T}\!\varphi^{\prime}(Y_{s})Z_{s}\cdot dW_{s},$
@@
$\tfrac{1}{2}\,\mathbb{E}\!\Big[\!\int_{\tau}^{T}\!|Z_{s}|^{2}ds\,\Big|\,\mathcal{F}_{\tau}\Big]$ $\leq\mathbb{E}\!\big[\varphi(Y_{T})-\varphi(Y_{\tau})\,\big|\,\mathcal{F}_{\tau}\big]+C_{Y}K(T-\tau)+\tfrac{3}{2}K_{\tilde{z}}C_{Y}\!\!\int_{0}^{T}\!\mathbb{E}|Z_{s}|^{2}ds$
@@
$\tfrac{1}{2}\,{\mathbb{E}}\!\Big[\!\int_{\tau}^{T}\!|Z_{s}|^{2}ds\,\Big|\,\mathcal{F}_{\tau}\Big]\leq(3K_{z})^{-1}e^{3K_{z}\|Y\|_{\mathbb{S}^{\infty}}}\left(\|\xi\|_{\infty}+K\,T+\tfrac{3}{2}K_{\tilde{z}}\!\!\int_{0}^{T}\!\!{\mathbb{E}}|Z_{s}|^{2}ds\right).$ (4.11)
@@
Step 2 (An explicit $\mathbb{H}^{2}$ bound). Write $\alpha:=K_{\tilde{z}}/K_{z}>0$. We may, without loss of generality (enlarge $K_{z}$ or $K_{\tilde{z}}$) assume $\alpha$ satisfies
$0<\alpha\;<\;\min\!\Big\{\tfrac{1}{2},\ \frac{e^{3K_{z}}-1}{e^{3K_{z}\|Y\|_{\mathbb{S}^{\infty}}}}\Big\}.$
Define positive constants
@@
Note $c_{1}$ can also be expressed as $c_{1}=\tfrac{1}{2}\big(\alpha e^{3K_{z}\|Y\|_{\mathbb{S}^{\infty}}}+1\big)=\frac{3K_{\tilde{z}}}{2}C_{Y}+\tfrac{1}{2}$. By the choice of $\alpha$, it is easy to verify $c_{2}>c_{1}$. By density of rational numbers, pick any $c\in(c_{1},c_{2})$ and consider the same $C^{2}$ function at in Step 1, so that $\varphi(0)=0$ and, for all $x$,
$|\varphi^{\prime}(x)|=\frac{e^{3K_{z}|x|}-2c}{3K_{z}}>\frac{e^{3K_{z}|x|}-2c_{2}}{3K_{z}}=0.$
So that we have identity $\varphi^{\prime\prime}(\cdot)-3K_{z}\left|\varphi^{\prime}(\cdot)\right|=2c$. Applying Itô to $\varphi(Y)$ on $[0,T]$ and arguing as in Step 1 gives
@@
By the definition of $c_{1}$ and $c>c_{1}$ we have $\frac{3K_{\tilde{z}}}{2}C_{Y}\leq c-\tfrac{1}{2}$. Hence
@@
Suppose Assumptions $(\rm H_{0})$} and $(\rm H_{q})$} hold, and that we work with the shifted generator satisfying
$g(t,0,\tilde{z})=0,\qquad\forall\,(t,\tilde{z})\in[0,T]\times\mathbb{R}^{d}$
(cf. Remark 2.4). Then for any solution $(Y,Z)\in\mathbb{S}^{2}\times\mathbb{H}^{2}_{\mathrm{BMO}}$ of BSDE (2.5), we have $Y\in\mathbb{S}^{\infty}$ and the following estimate holds:
@@
The proof relies on a linearization argument combined with the BMO property. Under the standing normalisation $g(t,0,\tilde{z})=0$, the BSDE (2.5) can be rewritten as
@@
$L_{s}:=\frac{g(s,Z_{s},\tilde{Z}_{s})-g(s,0,\tilde{Z}_{s})}{\left|Z_{s}\right|^{2}}\,Z_{s}\,\mathbf{1}_{\{\left|Z_{s}\right|>0\}},\qquad 0\leq s\leq T.$
@@
Since $Z\in\mathbb{H}^{2}_{\mathrm{BMO}}$ by assumption, $\big(\int_{0}^{t}\Big(\int_{\tilde{\Omega}}L_{s}\,d\tilde{\mathbb{P}}\Big)\cdot dW_{s}\big)_{0\leq t\leq T}$ is a BMO martingale, and its stochastic exponential $\mathcal{E}_{t}$ satisfies Kazamaki’s condition. Define the probability measure $\mathbb{P}^{L}$ by $d\mathbb{P}^{L}:=\mathcal{E}_{T}\,d\mathbb{P}$. By the above BMO argument, $(\mathcal{E}_{t})$ is a uniformly integrable martingale, so $\mathbb{P}^{L}$ is equivalent to $\mathbb{P}$. Girsanov’s theorem then implies that the process
@@
Since $\mathbb{P}^{L}$ is equivalent to $\mathbb{P}$, the same bound holds $\mathbb{P}$–almost surely (Lemma 4.3), and taking the essential supremum over $t\in[0,T]$ completes the proof. ∎
Then we have a direct result from previous two propositions.
@@
where $C$ is a constant depending only on $\left\|\xi\right\|_{L^{\infty}}$, $T$, and the constants appearing in Assumptions $(\rm H_{0})$} and $(\rm H_{q})$}.
@@
For $i=1,2$, let $(Y^{i},Z^{i})\in\mathbb{S}^{\infty}\times\mathbb{H}^{2}_{\mathrm{BMO}}$ solve
@@
with $\xi^{1},\xi^{2},g$ satisfying $(\mathrm{H}_{0})$ and $(\mathrm{H}_{g})$. Then
$\|Y^{1}-Y^{2}\|_{\mathbb{S}^{2}}^{2}+\|Z^{1}-Z^{2}\|_{\mathbb{H}^{2}}^{2}\ \leq\ C_{\mathrm{global}}\;\|\xi^{1}-\xi^{2}\|_{L^{4}},$
where $C_{\mathrm{global}}$ depends only on $T$, $K_{z}$, $K_{\tilde{z}}$ and $K_{0}$.
@@
Preliminary step (linearization and RHI). As in the proof of Proposition 4.4, we use a linearization argument, there exists processes $L_{s}$ and $\tilde{L}_{s}$ live on the product space such that, for all $s\in[0,T]$,
@@
and $\Delta g(s,Z_{s},\tilde{Z}_{s})=L_{s}\Delta Z_{s}+\tilde{L}_{s}\,\Delta\tilde{Z}_{s}$. Fix a small $\delta>0$ and write $I:=[T-\delta,T]$. According to Proposition 4.2, we have $Z^{1},Z^{2}\in\mathbb{H}^{2}_{\mathrm{BMO}}$, so in particular $\big(\int_{0}^{t}\tilde{\mathbb{E}}[L_{s}]\cdot dW_{s}\big)_{0\leq t\leq T}$ is a $\mathrm{BMO}$ martingale by (4.17). Moreover, by Corollary 4.5, there exists a constant $C>0$, depending only on $(T,K_{z},K_{\tilde{z}},K_{0})$, such that
@@
Hence $\Lambda:=\mathcal{E}(\int_{0}^{\cdot}\widetilde{\mathbb{E}}[L_{s}]\cdot dW_{s})$ defines an equivalent probability measure $\mathbb{P}^{L}$ by Girsanov’s theorem with Brownian motion $W^{L}_{t}:=W_{t}+\int_{0}^{t}\widetilde{\mathbb{E}}[L_{s}]\,ds$. In addition, by Lemma 2.2, the same processes admit $\mathrm{BMO}$ norms under the equivalent measure $\mathbb{P}^{L}$, up to multiplicative constants depending on the same constant. We then obtain,
$\Delta Y_{u}=\Delta\xi+\int_{u}^{T}\widetilde{\mathbb{E}}[\tilde{L}_{s}\,\Delta\tilde{Z}_{s}]\,ds-\int_{u}^{T}\Delta Z_{s}\cdot dW^{L}_{s},\qquad u\in I.$ (4.19)
The Doléans–Dade exponential $\Lambda=\mathcal{E}\!\left(\int_{0}^{\cdot}\tilde{\mathbb{E}}[L_{s}]\cdot dW_{s}\right)$ satisfies RHI from the Lemma 2.1,
$\bar{{\mathbb{E}}}\!\left[\left(\Lambda_{T}\right)^{q}\,\big|\,\bar{\mathcal{F}}_{t}\right]\ \leq\ C_{q}^{*}\,\left(\Lambda_{t}\right)^{q},\qquad 0\leq t\leq T,\quad 1\leq q\leq q^{*},$
@@
By the sliceability of BMO martingales and the function $\phi$ is strictly decreasing with respect to BMO norm, we can restrict to a short interval $I=[T-\delta,T]$ so that the local norm $\big\|\int_{0}^{\cdot}\tilde{\mathbb{E}}[L_{s}]\cdot dW_{s}\big\|_{\mathrm{BMO}(I)}$ is small enough to ensure $q^{*}>2$. Then we have the RHI for every nonnegative $X\in\mathcal{F}_{T}$, and we pick $q=2$:

### v2

@@ 4 A priori estimates and stability for BSDEs
In this section, we will work under the assumptions $(\rm H_{0})$} and $(\rm H_{q})$} to establish the stability of the mean-field BSDE (2.6), as stated in Theorem 4.6. The strategy of the proof is as follows. We first derive a priori upper bounds on the $\mathbb{S}^{\infty}$ norm of $Y$ and the BMO norm of $\int_{0}^{\cdot}Z_{s}\!\cdot dW_{s}$ for any solution $(Y,Z)$ to the BSDE (2.6). These estimates, together with fundamental properties of BMO martingales, lead to the stability result. We begin with a useful estimate that will play an important role in the analysis.
@@
Suppose that (H0) and (Hq) hold. Let $(Y,Z)\in\mathbb{S}^{\infty}\times\mathbb{H}^{2}$ be a solution to the mean-field BSDE (2.6). Then the stochastic process $\left(\int_{0}^{t}Z_{s}\!\cdot dW_{s}\right)_{0\leq t\leq T}$ defines a BMO martingale. Moreover, its BMO norm satisfies the following bound:
@@
The argument follows from arguments in Proposition 2.1 of briand2008stobsdes with modified $C^{2}$ function
$\varphi(x):=\frac{e^{3K_{z}|x|}-1-6cK_{z}|x|}{(3K_{z})^{2}},\qquad c\in\mathbb{R},x\in\mathbb{R}.$
@@
Step 1 (Estimate of $\|Z\|_{\mathbb{H}^{2}_{\mathrm{BMO}}}$ in terms of $\|Z\|_{\mathbb{H}^{2}}$). Fix $c=\tfrac{1}{2}$ then $\varphi(0)=0$ and, for all $x\in\mathbb{R}$,
$\varphi(x)\geq 0,\qquad\varphi^{\prime}(x)=\frac{e^{3K_{z}|x|}-1}{3K_{z}}\operatorname{sgn}(x),\qquad\varphi^{\prime\prime}(x)=e^{3K_{z}|x|},\qquad\varphi^{\prime\prime}(x)-3K_{z}\,|\varphi^{\prime}(x)|\equiv 1.$
@@
Let $\tau$ be any stopping time and apply Itô’s formula to $\varphi(Y)$ and using the growth estimate 4.9 for $g$ together with the symmetry $\widetilde{\mathbb{E}}|\tilde{Z}_{s}|^{2}=\mathbb{E}|Z_{s}|^{2}$ yields
$\varphi(Y_{\tau})$ $=\varphi(\xi)+\!\int_{\tau}^{T}\!\Big(-\varphi^{\prime}(Y_{s})\,\widetilde{\mathbb{E}}[g(s,Z_{s},\tilde{Z}_{s})]-\tfrac{1}{2}\,\varphi^{\prime\prime}(Y_{s})|Z_{s}|^{2}\Big)ds-\!\int_{\tau}^{T}\!\varphi^{\prime}(Y_{s})Z_{s}\cdot dW_{s}$
$\leq\varphi(\xi)+\!\int_{\tau}^{T}\!\Big(\frac{3K_{z}}{2}|\varphi^{\prime}(Y_{s})|-\tfrac{1}{2}\varphi^{\prime\prime}(Y_{s})\Big)|Z_{s}|^{2}ds+\!\int_{\tau}^{T}\!|\varphi^{\prime}(Y_{s})|\Big(K+\tfrac{3}{2}K_{\tilde{z}}\,\mathbb{E}|Z_{s}|^{2}\Big)ds$
$\leq\varphi(\xi)-\tfrac{1}{2}\!\int_{\tau}^{T}\!|Z_{s}|^{2}ds+\!\int_{\tau}^{T}\!|\varphi^{\prime}(Y_{s})|\Big(K+\tfrac{3}{2}K_{\tilde{z}}\,\mathbb{E}|Z_{s}|^{2}\Big)ds-\!\int_{\tau}^{T}\!\varphi^{\prime}(Y_{s})Z_{s}\cdot dW_{s},$
@@
$\tfrac{1}{2}\,\mathbb{E}\!\Big[\!\int_{\tau}^{T}\!|Z_{s}|^{2}ds\,\Big|\,\mathcal{F}_{\tau}\Big]$ $\leq\mathbb{E}\!\big[\varphi(\xi)-\varphi(Y_{\tau})\,\big|\,\mathcal{F}_{\tau}\big]+C_{Y}K(T-\tau)+\tfrac{3}{2}K_{\tilde{z}}C_{Y}\!\!\int_{0}^{T}\!\mathbb{E}|Z_{s}|^{2}ds$
@@
$\tfrac{1}{2}\,{\mathbb{E}}\!\Big[\!\int_{\tau}^{T}\!|Z_{s}|^{2}ds\,\Big|\,\mathcal{F}_{\tau}\Big]\leq C_{Y}\left(\|\xi\|_{L^{\infty}}+K\,T+\tfrac{3}{2}K_{\tilde{z}}\!\!\int_{0}^{T}\!\!{\mathbb{E}}|Z_{s}|^{2}ds\right).$ (4.12)
@@
Step 2 (An explicit $\mathbb{H}^{2}$ bound). Write $\alpha:=K_{\tilde{z}}/K_{z}>0$. We may, without loss of generality, assume $\alpha$ satisfies
$0<\alpha\;<\;\min\!\Big\{\tfrac{1}{2},\ \frac{e^{3K_{z}}-2}{e^{3K_{z}\|Y\|_{\mathbb{S}^{\infty}}}}\Big\}.$
by enlarging $K_{z}$ or $K_{\tilde{z}}$. Define positive constants
@@
Note that, by definition of $\alpha$, $c_{1}$ can also be expressed as $c_{1}=\tfrac{1}{2}\big(\alpha e^{3K_{z}\|Y\|_{\mathbb{S}^{\infty}}}+1\big)$. By the choice of $\alpha$, it is easy to verify $c_{1}<c_{2}$. Pick any $c\in(c_{1},c_{2})$ and consider the same $\varphi$ at in Step 1 with a different $c$ value, so that $\varphi(0)=0$ and, for all $x$,
$|\varphi^{\prime}(x)|=\frac{e^{3K_{z}|x|}-2c}{3K_{z}}>\frac{e^{3K_{z}|x|}-2c_{2}}{3K_{z}}\geq 0.$
We also have identity $\varphi^{\prime\prime}(\cdot)-3K_{z}\left|\varphi^{\prime}(\cdot)\right|=2c$. Applying Itô to $\varphi(Y)$ on $[0,T]$ and arguing as in Step 1 with $\tau=0$ gives
@@
By the definition of $c_{1}$ and the fact that $c>c_{1}=\frac{3K_{\tilde{z}}}{2}C_{Y}+\tfrac{1}{2}$ we have $\frac{3K_{\tilde{z}}}{2}C_{Y}<c-\tfrac{1}{2}$. Hence
@@
Suppose Assumptions $(\rm H_{0})$} and $(\rm H_{q})$} hold. Then for any solution $(Y,Z)\in\mathbb{S}^{2}\times\mathbb{H}^{2}_{\mathrm{BMO}}$ of BSDE (2.6), we have $Y\in\mathbb{S}^{\infty}$ and the following estimate holds:
@@
The proof relies on a linearization argument combined with the BMO property. Using Remark 2.6, without the loss of generality, we assume
Then, the BSDE (2.6) can be rewritten as
@@
$L_{s}:=\frac{g(s,Z_{s},\tilde{Z}_{s})-g(s,0,\tilde{Z}_{s})}{|\Delta Z_{s}|^{2}}\,\Delta Z_{s}\,\mathbf{1}_{\{\Delta Z_{s}\neq 0\}}\qquad 0\leq s\leq T,$
@@
Since $Z\in\mathbb{H}^{2}_{\mathrm{BMO}}$ by assumption, $\big(\int_{0}^{t}\Big(\int_{\tilde{\Omega}}L_{s}\,d\tilde{\mathbb{P}}\Big)\cdot dW_{s}\big)_{t\in[0,T]}$ is a BMO martingale, and its stochastic exponential $\mathcal{E}_{t}$ satisfies Kazamaki’s condition. Define the probability measure $\mathbb{P}^{L}$ by $d\mathbb{P}^{L}:=\mathcal{E}_{T}\,d\mathbb{P}$. By the above BMO argument, $(\mathcal{E}_{t})$ is a uniformly integrable martingale, so $\mathbb{P}^{L}$ is equivalent to $\mathbb{P}$. Girsanov’s theorem then implies that the process
@@
Since $\mathbb{P}^{L}$ is equivalent to $\mathbb{P}$, the same bound holds $\mathbb{P}$–almost surely by Lemma 4.3, and taking the essential supremum over $t\in[0,T]$ completes the proof. ∎
Then we have a direct result from the previous two propositions.
@@
where $C$ is a constant depending only on $(T,K_{z},K_{\tilde{z}},K_{0})$.
@@
For $i=1,2$, let $(Y^{i},Z^{i})\in\mathbb{S}^{\infty}\times\mathbb{S}^{\infty}$ solve
@@
with $\xi^{1},\xi^{2},g$ satisfying $(\mathrm{H}_{0})$ and $(\mathrm{H}_{q})$. Then there exist $p^{*}>1$ and, for every $p>p^{*}$, a constant $C_{p}>0$ such that
$\|Y^{1}-Y^{2}\|_{\mathbb{S}^{2p}}+\|Z^{1}-Z^{2}\|_{\mathbb{H}^{p}}\ \leq\ C_{p}\;\|\xi^{1}-\xi^{2}\|_{L^{2p}},$
where $p^{*}>1$ and $C_{p}$ depend only on $T$, $K_{z}$, $K_{\tilde{z}}$, $K_{0}$, and do not depend on $\|Z^{i}\|_{\mathbb{S}^{\infty}}$.
@@
Preliminary step (linearization, RHI and sliceability). As in the proof of Proposition 4.4, we use a linearization argument: there exist processes $L_{s}$ and $\tilde{L}_{s}$ on the product space such that, for all $s\in[0,T]$,
$L_{s}:=\frac{g(s,Z^{1}_{s},\tilde{Z}^{1}_{s})-g(s,Z^{2}_{s},\tilde{Z}^{1}_{s})}{|\Delta Z_{s}|^{2}}\,\Delta Z_{s}\,\mathbf{1}_{\{\Delta Z_{s}\neq 0\}},\qquad\tilde{L}_{s}:=\frac{g(s,Z^{2}_{s},\tilde{Z}^{1}_{s})-g(s,Z^{2}_{s},\tilde{Z}^{2}_{s})}{|\Delta\tilde{Z}_{s}|^{2}}\,\Delta\tilde{Z}_{s}\,\mathbf{1}_{\{\Delta\tilde{Z}_{s}\neq 0\}}$
with estimates
@@
and $\Delta g(s,Z_{s},\tilde{Z}_{s})=L_{s}\Delta Z_{s}+\tilde{L}_{s}\,\Delta\tilde{Z}_{s}$.
According to Proposition 4.2, we have $Z^{1},Z^{2}\in\mathbb{H}^{2}_{\mathrm{BMO}}$, so in particular $\big(\int_{0}^{t}\tilde{\mathbb{E}}[L_{s}]\cdot dW_{s}\big)_{t\in[0,T]}$ is a $\mathrm{BMO}$ martingale by (4.18) and symmetry. Moreover, by Corollary 4.5, there exists a constant $C>0$, depending only on $(T,K_{z},K_{\tilde{z}},K_{0})$, such that
@@
Hence $\Lambda:=\mathcal{E}(\int_{0}^{\cdot}\widetilde{\mathbb{E}}[L_{s}]\cdot dW_{s})$ defines an equivalent probability measure $\mathbb{P}^{L}$ by Girsanov’s theorem with Brownian motion $W^{L}_{t}:=W_{t}+\int_{0}^{t}\widetilde{\mathbb{E}}[L_{s}]\,ds$. In addition, by Lemma 2.2, the same processes admit $\mathrm{BMO}$ norms under the equivalent measure $\mathbb{P}^{L}$, up to multiplicative constants depending only on $C$. Under $\mathbb{P}^{L}$ we can rewrite (4.17) as
$\Delta Y_{u}=\Delta\xi+\int_{u}^{T}\widetilde{\mathbb{E}}[\tilde{L}_{s}\,\Delta\tilde{Z}_{s}]\,ds-\int_{u}^{T}\Delta Z_{s}\cdot dW^{L}_{s},\qquad u\in[0,T].$ (4.20)
The Doléans–Dade exponential $\Lambda=\mathcal{E}\!\left(\int_{0}^{\cdot}\tilde{\mathbb{E}}[L_{s}]\cdot dW_{s}\right)$ satisfies a reverse Hölder inequality by Lemma 2.1: there exists $q^{*}>1$ such that, for all $1\leq q\leq q^{*}$,
$\bar{{\mathbb{E}}}\!\left[\left(\Lambda_{T}\right)^{q}\,\big|\,\bar{\mathcal{F}}_{t}\right]\ \leq\ C_{q}^{*}\,\left(\Lambda_{t}\right)^{q},\qquad 0\leq t\leq T,$
@@
By boundedness of $\Big\|\int_{0}^{\cdot}\tilde{\mathbb{E}}[L_{s}]\cdot dW_{s}\Big\|_{\mathrm{BMO}}$, the constants $q^{*},C_{q}^{*}$ depend only on $(T,K_{z},K_{\tilde{z}},K_{0})$. Finally, set
$M_{t}:=\int_{0}^{t}(\tilde{\mathbb{E}}|K_{\tilde{z}}\big(1+|\tilde{Z}_{s}^{1}|+|\tilde{Z}_{s}^{2}|\big)|^{2})^{1/2}dW_{s},\qquad t\in[0,T].$
Then $M\in\mathcal{H}^{\infty}(\mathbb{P})$, and by schachermayer1996institut it is sliceable. Furthermore, since $Z^{1},Z^{2}\in\mathbb{S}^{\infty}$, we can (by (jackson2022triangle, Remark 2.8)) construct a deterministic partition $0=t_{0}<t_{1}<\dots<t_{N}=T$ with arbitrarily small mesh size bounded by a prescribed $\varepsilon>0$.
$\mathbb{S}^{p}$–control of $\Delta Y$. Fix $\varepsilon$, $k\in\{0,\dots,N-1\}$ and write $I_{k}=[t_{k},t_{k+1}]$. By subtracting (4.20) at $u=t_{k+1}$ from that at $u\in I_{k}$, we obtain the local BSDE
$\Delta Y_{u}=\Delta Y_{t_{k+1}}+\int_{u}^{t_{k+1}}\widetilde{\mathbb{E}}[\tilde{L}_{s}\,\Delta\tilde{Z}_{s}]\,ds-\int_{u}^{t_{k+1}}\Delta Z_{s}\cdot dW^{L}_{s},\qquad u\in I_{k}.$
Using Cauchy-Schwarz on the copy space, the growth bound (4.18), and then Fefferman’s inequality (Theorem 2.5 in kazamaki2006BMO) on $I_{k}$, we obtain
$\int_{u}^{t_{k+1}}\tilde{\mathbb{E}}|\tilde{L}_{s}||\Delta\tilde{Z}_{s}|ds$ $\leq\mathbb{E}\int_{0}^{T}(\tilde{\mathbb{E}}\big|K_{\tilde{z}}\big(1+|\tilde{Z}_{s}^{1}|+|\tilde{Z}_{s}^{2}|\big)\big|^{2})^{1/2}(\tilde{\mathbb{E}}|\Delta\tilde{Z}_{s}|^{2})^{1/2}ds$ (4.21)
$\leq\sqrt{2}\left\|\int_{0}^{\cdot}(\tilde{\mathbb{E}}|K_{\tilde{z}}\big(1+|\tilde{Z}_{s}^{1}|+|\tilde{Z}_{s}^{2}|\big)|^{2})^{1/2}dW_{s}\right\|_{BMO}\int_{u}^{t_{k+1}}(\mathbb{E}|\Delta Z_{s}|^{2})^{1/2}ds$
$\leq\varepsilon\int_{u}^{t_{k+1}}(\mathbb{E}|\Delta Z_{s}|^{2})^{1/2}ds\qquad u\in I_{k}.$
Taking $\mathbb{E}^{L}[\cdot\,|\,\mathcal{F}_{u}]$ in the local BSDE yields
$|\Delta Y_{u}|$ $\leq\mathbb{E}^{L}[|\Delta Y_{t_{k+1}}|\,|\,\mathcal{F}_{u}]+\varepsilon\int_{u}^{t_{k+1}}\big(\mathbb{E}|\Delta Z_{s}|^{2}\big)^{1/2}ds,\qquad u\in I_{k}.$
Iterating this estimate over the intervals $I_{k+1},I_{k+2},\dots,I_{N-1}$ and using $\Delta Y_{T}=\Delta\xi$, we obtain